OpenStudy (loser66):

Converge or diverge? \(\alpha >0, \alpha \in \mathbb R\) \[\sum_{n=1}^{\infty} sin (n^{-\alpha})\]

OpenStudy (rational):

recall that for \(x\gt 0\) we have \(\sin(x)\lt x\) from this we get \[\sin(n^{-\alpha}) \lt n^{-\alpha}\] since \(\sum\limits_{n=1}^{\infty}n^{-\alpha}\) converges for \(\alpha\gt 1\), the series \(\sum\limits_{n=1}^{\infty} \sin (n^{-\alpha})\) also converges for \(\alpha \gt 1\). we still need to see the behavior for \(0\lt\alpha\le 1\)

OpenStudy (loser66):

we have as t --> 0, sint / t --->1, can we use it?

OpenStudy (rational):

for testing the convergence in interval (0, 1] ?

OpenStudy (loser66):

yes. If we use -1 < sin t < 1 where t = 1/ n ^ alpha

OpenStudy (rational):

Ahh are you allowed to use limit comparison test ?

OpenStudy (rational):

\(a_n=\sin(\frac{1}{n^{\alpha}})\) \(b_n = \frac{1}{n^{\alpha}}\) \[\lim\limits_{n\to\infty}\dfrac{a_n}{b_n} =\lim\limits_{n\to\infty}\dfrac{\sin(\frac{1}{n^{\alpha}})}{\frac{1}{n^{\alpha}}}=1 \] therefore the series in question will have same behavior as p-series. that is, the series covnerges for \(\alpha\gt 1\) and diverges for \(\alpha\le 1\)

OpenStudy (loser66):

Yes,

OpenStudy (loser66):

Wow!! how easy it is to you. Thank you so much. I do appreciate.

OpenStudy (rational):

sinx/x=1 limit made it easy :) thanks to you too!

OpenStudy (loser66):

I need help understanding the proof. Please theorem 1.2 http://www.ux1.eiu.edu/~kparwani/integrals.pdf